5 years ago a person was 7 time as old as his son but at peresent he is only 4 time as old as his son how many time as old as his son,wil he be 5 years from now

Answer by lwsshak3(11628)

5 years ago a person was 7 time as old as his son but at peresent he is only 4 time as old as his son how many time as old as his son,wil he be 5 years from now.
***
5 years ago:
x=son's age
7x=father's age
..
at present:
(x+5)=son's age
(7x+5)=father's age
7x+5=4(x+5)
7x+5=4x+20
3x=15
x=5
son's age=x+5=10
father's age=7x+5=40
..
5 years from now:
son's age=10+5=15
father's age=40+5=45
father is how many times as old as his son=45/15=3